OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

@undeadknight

OpenStudy (anonymous):

@Jadeishere

OpenStudy (jadeishere):

Dude substitute the coordinates into the variables

OpenStudy (jadeishere):

when it gives you options it's easy

OpenStudy (anonymous):

howwwwww

OpenStudy (anonymous):

Just rewrite the problem

OpenStudy (texaschic101):

do you prefer elimination or substitution ?

OpenStudy (anonymous):

idk im pretty terrible at anything having to do with algebra

OpenStudy (anonymous):

so it would be -1, 3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Ok whats -2*-1+3

OpenStudy (texaschic101):

-2x + y = 1 --->(-1) -4x + y = -1 ------------- 2x - y = -1 (result of multiplying by -1) -4x + y = -1 -------------add -2x = -2 x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 1 + 2 y = 3 check.. -4x + y = -1 -4(1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct) solution is : (1,3)

OpenStudy (texaschic101):

any questions ?

OpenStudy (texaschic101):

that was solved by elimination...would you like me to show you how to solve it with substitution ?

OpenStudy (anonymous):

yea can ou help me with another

OpenStudy (texaschic101):

I can

OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (texaschic101):

substitution or elimination ?

OpenStudy (texaschic101):

I will do substitution, so you can see how both ways work...give me a sec

OpenStudy (anonymous):

it doesnt matter to me

OpenStudy (texaschic101):

x - 2y = 14 x = 2y + 14 now sub 2y + 14 in for x in the other equation x + 3y = 9 2y + 14 + 3y = 9 -- subtract 14 from both sides 2y + 3y = 9 - 14 -- combine like terms 5y = - 5 y = -1 now sub -1 in for y in either equation to find x x - 2y = 14 x - 2(-1) = 14 x + 2 = 14 x = 14 - 2 x = 12 check.. x + 3y = 9 12 + 3(-1) = 9 12 - 3 = 9 9 - 9 (correct) solution is : (12,-1)

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (texaschic101):

4x - 3y = 9 -8x + 6y = -18 reduces to -4x + 3y = -9 so we now have : 4x - 3y = 9 -4x + 3y = -9 ------------add 0 = 0 INFINITELY MANY SOLUTIONS -- answer is C

OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (texaschic101):

y = -5x + 9 now sub -5x + 9 in for y in the other equation 3x + 2y = 4 3x + 2(-5x + 9) = 4 3x -10x + 18 = 4 3x - 10x = 4 - 18 -7x = -14 x = 2 now sub 2 in for x in either equation to find y y = -5x + 9 y = -5(2) + 9 y = -10 + 9 y = -1 check.. 3x + 2y = 4 3(2) + 2(-1) = 4 6 - 2 = 4 4 = 4 (correct) solution is : (2,-1)